A man of mass 83 kg (weight 180 lb) jumps down to a concrete patio from a window ledge only 0.48 m (= 1.6 ft) above the ground. He neglects to bend his knees on landing, so that his motion is arrested in a distance of about 2.2 cm (= 0.87 in.). (a) What is the average acceleration  of the man from the time his feet first touch the  patio to the time he is brought fully to rest? (b) With what average force does this jump jar his bone structure?

Grade:10

1 Answers

Abhinav
8 years ago
for first part
2gh=-2ad
a = -gh/d
if we use g=10 m/s2
a= – 4.8*100 /2.2 = -218.18 m/s2

and
hence ,average force on man is = a*m